# Fluid dynamics, Statistical physics, Theoretical physics

### v2

@@ 4 Application I: Decaying Hydrodynamic Turbulence
where the averaging is over the solutions of the recurrent equation (31). The equally spaced points on a unit circle $\theta_{k}=\frac{2\pi k}{N}$ approximate the infinitesimal step $\theta\pm 0$ involved in the original loop equation. Such a function $f(\theta)$, with discontinuity at every angle, can exist only in the sense of distributions, which is why this limiting procedure is required. These subtle issues of discretization of the loop equation are discussed in the recent paper [Elia and Lellis(2025)], where they are justified under certain assumptions about the continuity of the loop and the flow.
The limit $N\to\infty$ should be taken in the end at fixed turbulent viscosity $\tilde{\nu}=\nu N^{2}$. The last relation was derived in [Migdal(2023)] and used later in [Migdal(2024h)]. It guarantees a finite limit for observable correlation functions, which was analytically computed in these papers, using QFT methods combined with number theory.
@@
The distribution of the variable $X(p,q)=\cot(\pi p/q)^{2}/N^{2}=\left(4R^{2}-1\right)/N^{2}$ for co-prime $1\leq p<q<N$ was computed in [Migdal(2024h)] by advanced number theory methods. In the statistical limit $N\to\infty$, this distribution remains discontinuous
@@
Figure 2: Verification of the Euler-ensemble theory against new high-resolution 4K DNS data [Sreenivasan and Rodhiya, Akash(2026)]. (Top Left): The predicted parabolic relation between time $t$ and the integral length scale $L_{M}(t)$. The simulation data is in excellent agreement with a parabolic fit. (Top Right): The parameter-free prediction for the effective index $f(x)=\langle r\partial_{r}\log(\Delta v^{2})\rangle$. The theory captures the universal curvature seen in the DNS data. (Bottom): Verification of the full energy decay law at $Re_{\lambda}=145$. The DNS data (blue line) deviates from the leading-order term of the theory ($E\propto L_{M}^{-5/2}$, gray dash-dotted line). However, it is in excellent agreement with the complete theoretical solution (orange line), which includes all sub-leading exponents from the Mellin-Barnes integral. Only the two unknown scales (energy and length) were fitted to match the DNS data with the theory; no dimensionless parameters were changed. Such a perfect fit with only two parameters for about a thousand data points provides a high level of confidence.
@@ 4.1 Key Prediction 1: The Energy Spectrum and Intermittency
The Euler ensemble provides a complete, parameter-free prediction for all statistical properties of decaying turbulence. Most notably, it determines the velocity correlation functions and the energy spectrum [Migdal(2024h)]. The computations are rather heavy; in addition to above distribution of the radius $R$ of the regular star polygons, they involve some QFT technology, such as path integrals in the WKB limit, related to the classical trajectory for the Ising spin density $\alpha(k/N)$ and the functional determinant for harmonic fluctuations around this trajectory, computed by the zeta regularization method.
@@
The function $V(p)$, derived from the theory (see Appendix C), has poles that define the universal intermittency exponents, $p_{i}$. This spectrum of indices is remarkable: it contains not only rational numbers but also an infinite series of complex-conjugate pairs, $p_{n}=7\pm it_{n}$, where $\frac{1}{2}\pm it_{n}$ are the nontrivial zeros of the Riemann zeta function. This prediction [Migdal(2024h)], which is in excellent agreement with high-resolution direct numerical simulations (DNS), establishes a deep and unexpected connection between turbulence and number theory.
A direct test of this predicted functional form is the effective index, $f(x)=\langle r\partial_{r}\log(\Delta v^{2})\rangle$, where $x=r/L(t)$. This test is shown in Figure 2 (Top Right), which compares the parameter-free theoretical curve (orange line) to the latest 4K DNS data (blue line) [Sreenivasan and Rodhiya, Akash(2026)]. The agreement is excellent across the primary range.
@@ 4.2 Key Prediction 2: The Geometry of Mixing
The Euler ensemble also dictates the behavior of a passive scalar (e.g., temperature or a dye) advected by the turbulent flow [Migdal(2025e)]. For a localized initial condition, the scalar does not spread into a simple Gaussian cloud. We derived the loop equation for the evolution of the passive scalar $T(r,t)$ from conventional diffusion-advection equations, coupled with the loop equations for the velocity circulation. The resulting solution for the scalar density in the extreme turbulent limit is quite unexpected. Instead of a smooth Gaussian cloud, it forms a series of expanding, quantized concentric shells (see Figure 3. The radial profile of the scalar concentration is piecewise-parabolic, and the shell structure is organized by the Euler totient function, which counts the number of valid star polygons of a given complexity. This sharp, geometric prediction offers a clear target for future experiments.
@@ 5 Application II: Magnetohydrodynamic (MHD) Turbulence
The loop space solution can be generalized to more complex systems, such as magnetohydrodynamics (MHD), which describes the dynamics of conducting fluids, such as astrophysical plasmas [Migdal(2025d)]. The system is now described by two coupled loop functionals, one for the velocity circulation ($\Gamma_{v}$) and one for the magnetic vector potential circulation ($\Gamma_{a}$), each governed by an interacting Euler ensemble.
@@ 6 The Yang-Mills Connection and Universal Duality
The final step reveals the framework’s profound universality. The Yang-Mills gradient flow, a central tool for understanding the vacuum structure and confinement problem in QCD, is governed by a non-Abelian version of the same loop dynamics [Migdal(2025c)]. Applying our framework, we find exact solutions for the evolution of the Wilson loop.
@@
and minimizes the dual area functional under this constraint. Its area $S[C]$ exactly solves the non-perturbative fixed-point Yang–Mills loop equation,
$\partial_{\mu}\frac{\delta S[C]}{\delta\sigma_{\mu\nu}(\theta)}=0,$ (49)
@@
$W[C]=\exp\left(-\kappa(S_{+}[C]+S_{-}[C])\right),$ (50)
where $S_{\pm}[C]$ are the areas of the analytic self-dual and anti-self-dual surfaces bounded by the same loop.
For planar loops, the dual area is $|S|=2\sqrt{2}|D|$, reproducing the confining area law, while for general nonplanar or twisted loops, the shape dependence distinguishes it from the Euclidean minimal surface. The harmonic-map solution [Migdal(2025a)] clarifies the geometry: the Hodge-dual surface corresponds to a harmonic embedding $X:\mathcal{D}\!\to\!M_{N}(\mathbb{C})$ with the induced conformal metric $g_{ab}=e^{\varphi}\delta_{ab}$, where the Liouville field $\varphi$ governs local scale fluctuations. This analytic result provides the first explicit geometric realization of the Yang–Mills confining string and establishes a bridge between gauge theory, dual string geometry, and the loop-space formulation of turbulence.

### v3

@@ 4 Application I: Decaying Hydrodynamic Turbulence
where the averaging is over the solutions of the recurrent equation (33). The equally spaced points on a unit circle $\theta_{k}=\frac{2\pi k}{N}$ approximate the infinitesimal step $\theta\pm 0$ involved in the original loop equation. Such a function $f(\theta)$, with discontinuity at every angle, can exist only in the sense of distributions, which is why this limiting procedure is required. These subtle issues of discretization of the loop equation are discussed in the recent paper [DeLellisInprep], where they are justified under certain assumptions about the continuity of the loop and the flow.
The limit $N\to\infty$ should be taken in the end at fixed turbulent viscosity $\tilde{\nu}=\nu N^{2}$. The last relation was derived in [migdal2023exact] and used later in [migdal2024quantum]. It guarantees a finite limit for observable correlation functions, which was analytically computed in these papers, using QFT methods combined with number theory.
Remark 8.
The requirement of fixed turbulent viscosity arises from the necessity that the energy dissipation rate in the infinite system
$\mathcal{E}=\frac{\nu\int_{V}\omega^{2}}{V}$ (35)
stays finite in the turbulent limit $\nu\to 0$, taken simultaneously with the local limit $N\to\infty$ when the approximating polygon becomes a continuous loop $C$. In this limit, our discrete MLE tends to the solution of the continuum loop equation. The scaling relation $\nu\propto 1/N^{2}$ follows from the estimate [migdal2023exact, migdal2024quantum] using the second area derivative to compute $\omega^{2}$ and shrinking the loop to a point. In this limit, the statistical average of (28) is determined by the WKB asymptotics, and this leads to dramatic simplification, eventually producing exact finite result for the vorticity correlator. This result involves the average over an ensemble of $\mathbb{Q}$ rational numbers, eventually leading to beautiful Mellin-Barnes integrals involving ratios of Riemann zeta functions [migdal2024quantum],section IX.
@@
The distribution of the variable $X(p,q)=\cot(\pi p/q)^{2}/N^{2}=\left(4R^{2}-1\right)/N^{2}$ for co-prime $1\leq p<q<N$ was computed in [migdal2024quantum] by advanced number theory methods. In the statistical limit $N\to\infty$, this distribution remains discontinuous
@@
Figure 2: Verification of the Euler-ensemble theory against new high-resolution $4096^{3}$ DNS data [SreeniAkash2025]. (Top Left): The predicted parabolic relation between time $t$ and the integral length scale $L_{M}(t)$. The simulation data is in excellent agreement with a parabolic fit. (Top Right): The parameter-free prediction for the effective index $f(x)=\langle r\partial_{r}\log(\Delta v^{2})\rangle$. The theory captures the universal curvature seen in the DNS data. (Bottom): Verification of the full energy decay law at $Re_{\lambda}=145$. The DNS data (blue line) deviates from the leading-order term of the theory ($E\propto L_{M}^{-5/2}$). However, it is in excellent agreement with the complete theoretical solution (orange line), which includes all sub-leading exponents from the Mellin-Barnes integral. Only the two unknown scales (energy and length) were fitted to match the DNS data with the theory; no dimensionless parameters were changed. Such a perfect fit with only two parameters for about a thousand data points provides a high level of confidence.
@@ 4.1 Key Prediction 1: The Energy Spectrum and Intermittency
The Euler ensemble provides a complete, parameter-free prediction for all statistical properties of decaying turbulence. Most notably, it determines the velocity correlation functions and the energy spectrum [migdal2024quantum]. The computations are rather heavy; in addition to above distribution of the radius $R$ of the regular star polygons, they involve some QFT technology, such as path integrals in the WKB limit, related to the classical trajectory for the Ising spin density $\alpha(k/N)$ and the functional determinant for harmonic fluctuations around this trajectory, computed by the zeta regularization method.
@@
The function $V(p)$, derived from the theory (see Appendix C), has poles that define the universal intermittency exponents, $p_{i}$. This spectrum of indices is remarkable: it contains not only rational numbers but also an infinite series of complex-conjugate pairs, $p_{n}=7\pm it_{n}$, where $\frac{1}{2}\pm it_{n}$ are the nontrivial zeros of the Riemann zeta function. This prediction [migdal2024quantum], which is in excellent agreement with high-resolution direct numerical simulations (DNS), establishes a deep and unexpected connection between turbulence and number theory.
A direct test of this predicted functional form is the effective index, $f(x)=\langle r\partial_{r}\log(\Delta v^{2})\rangle$, where $x=r/L(t)$. This test is shown in Figure 2 (Top Right), which compares the parameter-free theoretical curve (orange line) to the latest $4096^{3}$ DNS data (blue line) [SreeniAkash2025]. The agreement is excellent across the primary range.
@@
Remark 9.
It may be worth noting that this superposition of leading plus subleading power terms
$E(L)\to AL^{-5/2}+BL^{-11/2}+\dots$
will not imitate a mixture of powers. Asymptotically, the effective index
$n(L)=LE^{\prime}(L)/E(L)\to-\frac{5}{2}-\frac{3B}{AL^{3}}+\dots$
will rather look like a leading index plus a decreasing powerlike correction, so this will not be a rotated line in a log-log scale but rather a straight line with a leading slope, curved at the left end by a $1/L^{3}$ correction, just like it appears in Figure 2.
@@
LKB regime violation of universality as a boundary effect
The Euler ensemble solution applies to decaying turbulence in an infinite system at an infinite Reynolds number. Still, as we have seen, the DNS for a finite grid with $L=4096$ steps and Reynolds number $145$ already matches our prediction within statistical errors, up to the boundary corrections, showing that at large distances $r\sim L$ or small wavelengths $k\sim\pi/L$.
These boundary corrections are also related to the initial shape of the energy spectrum. The time decay of the energy spectrum starts at large $k$, where the K41 spectrum $E(k,0)\sim k^{-5/3}$ was imposed as initial data for lack of a better guess.
This spectrum decays at the high end, curves on a log scale, and effectively shifts all the points to the left. The lower part of the spectrum remains at its initial state at $t=0$. There are two popular choices: Birkhoff-Saffman regime (BS): $E(k,0)\propto k^{2}$ and Loitsianskii-Kolmogorov-Batchelor (LKB) regime $E(k,0)\propto k^{4}$. The BS regime is the most general; it requires no parameter tuning, and it is apparently observed in physical experiments on decaying turbulence in water tanks, pipes, and wind tunnels. The experimental data, corrected for the boundary effects, perfectly match our predictions $E_{tot}(t)\propto t^{-5/4}$.
In the DNS of the $4096^{3}$ grid that we analyzed above, the BS regime also fits our theory very well. However, as observed in [SreeniAkash2025], with the LKB initial data, the fit is not as good; there are systematic deviations that the DNS authors fit to the LKB version of the K41 law.
$E_{tot}(t)\approx t^{-10/7}.$
On the other hand, the data for the second-order velocity moment $\langle(\Delta\boldsymbol{v})^{2}\rangle(r,t)$ in the LKB regime [SreeniAkash2025] matched our theory as well as they did in the BS regime. Both observables being mathematical transformations of the same microscopic quantity—the decaying energy spectrum—this paradoxical discrepancy needs an explanation.
This explanation is simple: the effective index is the ratio of two integrals over the energy spectrum, with kernels supported on the bulk of the flow ($\pi/L\ll k\ll\pi/2$). This region is supposed to be described by our turbulent attractor– the Euler ensemble.
The total energy, as defined, equals the integral over the full spectrum. In the lattice system, this will be the sum for all quantized wavevectors $k=\pi n/L$. The first several terms in this sum are influenced by the initial $k^{2}$ or $k^{4}$ spectrum.
We do not expect our theory to match the low end of the energy spectrum: there are systematic errors of the lattice approximation to the true isotropic homogeneous turbulence. However, the theoretical spectrum grows at small $k$, so the total energy does not converge in the bulk. We had to cut off the spectral integral at some finite $k=k_{0}\sim 1/L$ in [migdal2024quantum] to predict the scaling law $E_{tot}(t)\sim t^{-5/4}$. The coefficient in front of this power law is not universal: it is proportional to the corresponding power of $k_{0}$.
To avoid these apparent violations of universality due to contamination from the initial/boundary data, we switch to the enstrophy, which is proportional to the time derivative of energy, so it can serve equally well to describe the energy decay
$E^{\prime}_{tot}(t)=-\nu En(t);$ (50)
$En(t)=\int_{0}^{\infty}dkk^{2}E(k,t);$ (51)
The scaling law for enstrophy in our theory is
$En(t)\propto(t+t_{0})^{-9/4}\propto L(t)^{-9/2}$ (52)
The $10/7$ law for the energy would translate into
$\text{K41: }En(t)\propto(t+t_{0})^{-17/7}\propto L(t)^{-34/7}$ (53)
The difference with the energy is that the enstrophy integral is dominated by the bulk, where the theory of turbulence would apply. Neither the boundary conditions nor the low end of the spectrum, contaminated by initial data, will contribute to the enstrophy integral. Moreover, the enstrophy, the mean square of the vorticity, does not depend on the potential part of the velocity field.
It is this potential component that is most affected by the initial data for the BS or LKB regimes. The coefficient in front of $k^{2}$ is proportional to the square of mean velocity, and that in front of $k^{4}$ - to the square of mean angular momentum. In general, the $k^{2}$ term dominates, but if one tunes parameters so that there is no mean velocity, the $k^{4}$ term is left as the first term of expansion at small $k$. Obviously, neither of these terms has anything to do with the turbulent regime; these are just initial data for the potential part of the velocity field. These initial macroscopic motions are related to the dynamics of the vortex structures responsible for turbulence through the advection they induce. Thus, the solution for the velocity field will initially differ in these two regimes (BS and LKB), but we expect the vortex dynamics to eventually settle on the turbulent attractor described by the Euler ensemble.
The DNS data can be safely summed without filtering out the lattice artifacts at $k\sim 1/L$. The integral will converge in the bulk, where the turbulence theory applies. Thus, we can use the entire dataset in the turbulent range where statistical equilibrium is reached, but the turbulent kinetic energy has not yet dissipated. That will give us better statistics than for the bulk energy decay. The enstrophy test is therefore a direct probe of the turbulent attractor, uncontaminated by boundary or initial‑condition effects.
We measured the enstrophy decay and compared it with the predicted scaling law. This comparison is made at Figs. 3.
Figure 3: The total enstrophy decay for the LKB regime, when $E(k,0)\propto k^{4}$. The four plots correspond to four different simulations with randomized initial data. The log of enstrophy $\log En(t)$ is plotted against $\log L(t)$ where $L(t)$ is defined as ratio of first and second moments of $k$ over the whole range of $k$ for a given time in the turbulent region of time moments $100<t<500$, where $t(L)$ is well approximated by a parabola. The upper limit 500500 was close to the end of available data for the LKB regime. The red dotted line – the DNS data from [SreeniAkash2025], the blue dashed line is our scaling law $En(t)\propto L^{-9/2}$, and the black dotted line corresponds to the K41 law $En(t)\propto L^{-34/7}$. The $9/2$ law fits the data perfectly, but the $34/7$ is not even close.
This remarkable verification of our scaling law for enstrophy (which is proportional to the time derivative of the energy) suggests that the ostensible violations of universality in the LKB regime for the energy decay were due to contributions from the initial and boundary data. Once we eliminated these contributions by an extra power of $k^{2}$ which is present in the enstrophy spectral integral, the universal scaling law of the Euler ensemble was restored. This resolves the issue of universality violation in the LKB regime as an artifact of the boundary conditions.
Remark 10.
Let us clarify the critical issue regarding the universality of the turbulent attractor described by the Euler ensemble and the violation of this universality by initial and boundary data.
In an infinite system with infinite initial energy, it requires infinite time to dissipate this energy. This scenario corresponds to ideal decaying turbulence, which is described exactly by the Euler ensemble in the limit of zero viscosity at a fixed energy dissipation rate (infinite Reynolds number). This constitutes the mathematical solution to the problem of decaying turbulence.
However, there is a distinct engineering problem: turbulent flow in a finite system of a particular shape, with finite energy and a finite initial Reynolds number. In such a system, the ideal Euler solution is modified by initial and boundary corrections. The bulk of the decaying energy spectrum in this finite system evolves toward the Euler ensemble solution. This explains why the effective index $\xi_{2}(r/\sqrt{t})$ in the DNS matches the Euler ensemble prediction perfectly within statistical error. The higher moments of velocity are also calculable in Euler ensemble, though it will take some effort (a good project for an ambitious postdoc, trained in QFT and familiar with mathematical physics and number theory).
Conversely, the boundary and initial corrections are related to the lower part of the spectrum, which grows into the region of small $k$ within the Euler ensemble solution. In a finite system, this growth is cut off at $k_{0}\sim 1/L$, where $L$ is the system size. This infrared part of the spectrum is not universal; in Direct Numerical Simulations (DNS), it is strongly influenced by lattice artifacts and finite-size effects. Here, the microscopic theory must step aside, leaving the arena to phenomenological engineering theories such as the K41 scaling laws. These laws reflect dimensional counting and thus apply regardless of the specific dynamics.
Nevertheless, as the box size increases, these boundary corrections eventually diminish. In the DNS [SreeniAkash2025], these corrections to the energy decay shift the effective exponents by a few percent. In physical experiments, which operate at much higher Reynolds numbers and in significantly larger domains, the decay index fits our theoretical prediction of $5/4$ perfectly.
In conclusion, the DNS data [SreeniAkash2025] provide robust validation of our theoretical predictions, agreeing within the statistical and systematic precision (a few percent) of $4096^{3}$ lattice simulations at Reynolds numbers $\sim 10^{2}$. This confirmation encompasses the universality of the decay scaling law $E^{\prime}(t)\propto t^{-9/4}$, the diffusive growth of the effective length $L(t)\propto\sqrt{t}$, and the specific behavior of the effective index (the logarithmic derivative of the second velocity moment as a function of $r/L(t)$). In contrast, alternative phenomenological scaling laws (K41, LKB, BS) fall outside the statistical confidence limits and are thus ruled out. Looking ahead, supercomputing capabilities are poised to surpass $32K^{3}$ grid resolutions, while wind tunnel experiments are already approaching Reynolds numbers in the tens of thousands. These next-generation studies will be decisive, providing the resolution necessary to disentangle boundary effects from bulk dynamics and to definitively verify the predicted universal attractor.
@@ 4.2 Key Prediction 2: The Geometry of Mixing
The Euler ensemble also dictates the behavior of a passive scalar (e.g., temperature or a dye) advected by the turbulent flow [migdal2025mixing]. For a localized initial condition, the scalar does not spread into a simple Gaussian cloud. We derived the loop equation for the evolution of the passive scalar $T(r,t)$ from conventional diffusion-advection equations, coupled with the loop equations for the velocity circulation. The resulting solution for the scalar density in the extreme turbulent limit is quite unexpected. Instead of a smooth Gaussian cloud, it forms a series of expanding, quantized concentric shells (see Figure 4). The radial profile of the scalar concentration is piecewise-parabolic, and the shell structure is organized by the Euler totient function, which counts the number of valid star polygons of a given complexity. This sharp, geometric prediction offers a clear target for future experiments.
Remark 11.
The velocity singularities vs those of passive scalar Let us clarify the difference between these strong singularities (concentric shells of discontinuities in the passive scalar mean density, condensing to a delta-like singularity in the extreme turbulent limit), and weaker singularities the Euler ensemble predicts in velocity correlations. For the velocity field, the two-point correlation function has a first singular term $\delta v^{2}\sim r^{2}+|r|^{5/2}$, which is weaker than $|r|^{2/3}$ of K41 model. As for the passive scalar, there are actual infinities and discontinuities, generated by the velocity field distributed by the same Euler ensemble. These singularities may be observed (in some smear form) as the above ramp-cliffs, but they are clealy impossible to observe in DNS on any finite grid. As we discuss in the research paper submitted to the same issue, the passive scalar accumulated in a spherical ball around initial injection point, maniveste less singular, but also quite distinctive behavior, which can be observed in DNS. Also, the Fourier transform of the passive scalar density is a smooth universal function of $k\sqrt{t}$, also measurable in DNS.
@@ 5 Application II: Magnetohydrodynamic (MHD) Turbulence
The loop space solution can be generalized to more complex systems, such as magnetohydrodynamics (MHD), which describes the dynamics of conducting fluids, such as astrophysical plasmas [migdal2025mhd]. The system is now described by two coupled loop functionals, one for the velocity circulation ($\Gamma_{v}$) and one for the magnetic vector potential circulation ($\Gamma_{a}$), each governed by an interacting Euler ensemble.
@@ 6 The Yang-Mills Connection and Universal Duality
The final step reveals the framework’s profound universality. The Yang-Mills gradient flow, a central tool for understanding the vacuum structure and confinement problem in QCD, is governed by a non-Abelian version of the same loop dynamics [migdal2025YangMills]. Applying our framework, we find exact solutions for the evolution of the Wilson loop.
@@
and minimizes the dual area functional under this constraint. The surface area is defined as
$S_{\chi}[C]=\min_{X}\int_{\Sigma}d^{2}\xi\sqrt{\Sigma_{\mu\nu}^{2}/2};$ (56)
The area element is defined as
$\Sigma_{\mu\nu}=\epsilon_{ab}\partial_{a}X^{i}_{\mu}\partial_{b}X^{i}_{\nu};$ (57)
The Hodge chirality $\chi=\pm 1$ enters through the boundary conditions
$X^{i}_{\mu}(\partial\Sigma)=\eta^{\chi,i}_{\mu\nu}C_{\nu};$ (58)
Here $\eta^{\chi,i}_{\mu\nu}$ are ’t Hooft’s matrices corresponding to Hodge duality $\chi=\pm 1$
$\eta^{\chi,i}_{\mu\nu}=\delta_{i\mu}\delta_{\nu 4}-\delta_{i\nu}\delta_{\mu 4}+\chi e_{i\mu\nu 4};$ (59)
Its area $S[C]$ is an exact zero more of the loop diffusion operator in four dimesions
$\mathcal{L}S[C]=0;$ (60)
And as a consequence,there is a symmetry in the non-perturbative fixed-point Yang–Mills loop equation (so called Makeenko-Migdal loop equation [MM1981NPB, Mig83]),
$\mathcal{L}(W[C])=\lambda\int_{C}dx_{\mu}\int_{C}dy_{\mu}\delta^{4}(x-y)W[C_{xy}]W[C_{yx}],$ (61)
@@
$W[C]=W_{0}[C]\exp\left(-\kappa(S_{+}[C]+S_{-}[C])\right),$ (62)
where $S_{\pm}[C]$ are the areas of the analytic self-dual and anti-self-dual surfaces bounded by the same loop, and $W_{0}[C]$ is the perturbative solution corresponding to asymptotic freedom. The minimality conditions with the extra self-duality constraint allow an exact harmonic solution inside the unit disk ($|z|<1$)
$X^{a}_{\mu}(z,\bar{z})=\eta^{\chi,i}_{\mu\nu}(f_{\nu}(z)+\bar{f}_{\nu}(\bar{z}));$ (63)
$C_{\mu}(\theta)=2\textbf{Re }f_{\nu}\left(e^{\mathrm{i}\theta}\right)$ (64)
For non-intersecting loops, the dual area is $S_{\pm}[C]=2\sqrt{2}|A[C]|$, where $A[C]$ is the minimal Euclidean area bounded by $C$. This reproduces the confining area law, while for self-intersecting loops, our solution is additive across the closed parts, unlike the Euclidean minimal surface, which switches between topologies to achieve the global minimum of the area. Our additive solution, however, represents the local minimum of the area, which is sufficient for the confining factor to satisfy the MM loop equations.
This analytic result provides the first explicit geometric realization of the Yang–Mills confining string and establishes a bridge between gauge theory, dual string geometry, and the loop-space formulation of turbulence.